Word problems get me so confused. Can someone point me in the right direction with these.
In a town election, the winning candidate had 220 more votes than the loser. If 810 votes were cast in all, how many votes did each candidate receive?

The length of a rectangle is 2 in. more than twice its width. If the perimeter of the rectangle is 34 in., find the dimensions of the rectangle.

A small firm produces both AM and AM/FM car radios. The AM radios take 15 h to produce, and the AM/FM radios take 20 h. The number of production hours is limited to 300 h per week. The plant’s capacity is limited to a total of 18 radios per week, and existing orders require that at least 4 AM radios and at least 3 AM/FM radios be produced per week. Write a system of inequalities representing this situation. Then, draw a graph of the feasible region given these conditions, in which x is the number of AM radios and y the number of AM/FM radios.

Answer by checkley75(3666)

x+(220+x)=810
2x=+220=810
2x=810-220
2x=590
x=590/2
x=295 votes for the looser & 295+220=515 for the winner.
proof
295+515=810
810=810
---------------------------------------------------------------
2(2W+2)+2W=34
4W+4+2W=34
6W=34-4
6W=30
W=30/6
W=5 FOR THE WIDTH.
L=2*5+2
L=10+2
L=12 FOR THE LENGTH.
PROOF
2*5=2*12=34
10+24=34
34=34